OpenStudy (anonymous):

Which pair of ratios forms a proportion? 10 over 22. and 5 over 44. 16 over 9. and 8 over 18. 7 over 8. and 14 over 16. 20 over 24. and 40 over 12.

OpenStudy (texaschic101):

ratios are nothing but equivalent fractions so is : 10/22 and 5/44 equivalent ? 10/22 = 5/11 and 5/44 = 1/11...so these are not equivalent, and thus do not form a proportion. 16/9 and 8/18.....8/18 = 4/9....these are not equivalent 7/8 and 14/16.....14/16 reduces to 7/8....these are equivalent fractions, thus, they make a proportion you can even check, like this 7/8 = 14/16 cross multiply (8)(14) = (7)(16) 112 = 112 (correct) --- this one is your answer understand ?